OpenStudy (anonymous):

Kara has 21 collectible buttons to sell. She sells the small buttons for $3 and the large buttons for $4, and earns a total of $68. If the small buttons are represented by x and the large buttons are represented by y, which of the following systems of equations can be used to calculate the number of large and small buttons sold?

OpenStudy (anonymous):

I am really unsure as to how I am supposed to do this.

OpenStudy (anonymous):

x + y = 21, 3x + 4y = 68 x + y = 68, 3x + 4y = 21 x + y = 21, 4x + 3y = 68 x + y = 68, 4x + 3y = 21

OpenStudy (anonymous):

Those are the answer choices.